70% of people wear seat belts. If 20 people were randomly checked, find the probability that 9 or fewer were wearing seat belts. (Note: This is a binomial problem since people either wear seat belts or they don't.)
Answer #1

P[ Person wearing seat belt ] = p = 70% = 0.7

The population proportion of success is p = 0.7, also, 1 - p = 1 - 0.7 = 0.3, and the sample size is n= 20. We need to compute

Therefore, we get that

This implies that
Answer #1

P[ Person wearing seat belt ] = p = 70% = 0.7

The population proportion of success is p = 0.7, also, 1 - p = 1 - 0.7 = 0.3, and the sample size is n= 20. We need to compute

Therefore, we get that

This implies that